“Hello World!” — α = Γ − \sqrt{x^{2} + y^{2}}.

The difficult first flight afflicts the efflux differently.

naïve; belovèd; protégé; rôle coördinator; soufflés,

crêpes, and pâtés; Æsop’s Œuvres, Ångstrom, Straße.

As far as TEX is concerned, CM fonts in ATM compatible Adobe Type 1 format (‘PostScipt’ or ‘ATM’ fonts ) are the same as CM fonts in bitmapped PK font format, since TEX itself only cares about TFM meric files. Consequently, there are no changes needed in your TEX source file when you use them.

The Type 1 versions of the CM fonts do have the ability to easily scale (in addition to providing different fonts for different design sizes). This means you can use ‘in-between’ sizes like CMR17 at 15pt. David Carlisle’s Type1CM package takes advantage of this. It may be found at:

/tex-archive/macros/latex/contrib/supported/type1cm on CTAN.

Read ‘type1cm.txt’ in that directory for details.

CM ‘text’ fonts have fixed encoding (character layout) called ‘TeX text’ or OT1. L^{A}TEX 2e uses OT1 encoding unless you tell it otherwise.

NOTE: the setting of ‘Font > Encoding’ in DVIWindo does not affect CM fonts since CM ‘text’ fonts have to be treated as pi/symbol/decorative/math fonts in order to prevent reencoding to platform specific layout (Macintosh standard roman or Windows ANSI).

If you need hyphenation in languages using accented and special characters consider the European Modern font set which has ready-made accented chars. The basic characters in European Modern are the same as those in CM. L^{A}TEX 2e support for EM fonts may be found in the ‘em’ package.